OpenStudy (anonymous):

Write the equation of a circle where A(1,−4)and B(−7,8)are the endpoints of a diameter.

OpenStudy (aum):

Find the midpoint of the two endpoints. That will be the center since they are endpoints of a diameter. Find the distance between the center and either one of the two given points. That will be the radius. Equation of the circle is: (x - h)^2 + (y - k)^2 = r^2 where (h,k) is the center and r is the radius of the circle.

OpenStudy (anonymous):

Thank you.. I tried working this out but i still getting the wrong answer.. could you show me how to plug everything into the formulas and the steps that you take? please? @aum

OpenStudy (anonymous):

Nevermind, I figured it out. Thank you so much!

OpenStudy (aum):

Okay.

OpenStudy (aum):

What equation did you get?

OpenStudy (anonymous):

(x+3)^2+(y-2)^2=52 I believe that is right.. at least it looks that way.

OpenStudy (aum):

You did good! :) (Sorry the notifications are not working properly and I did not know about your reply until long after it was posted.)